### v1

# Infrared spectra of methane-containing ice mixtures for JWST data analysis

Abstract
Context. Solid methane (CH4) is an important molecule in interstellar and planetary environments, serving as a precursor to complex organic compounds and a potential biosignature in exoplanetary studies. Despite its significance, laboratory data on low-temperature phase of methane below 10 K remain limited.
Aims. We aim to obtain spectra of methane in binary mixtures at 10 K and compare it to the spectra obtained at 6.7 K. These temperatures correspond to phases II and II* of pure methane and are representative of dark molecular clouds and protostars at early stages. We also aim to test the obtained data applicability to JWST data interpretation.
Methods. Laboratory reference spectra were obtained on the ISEAge setup via FTIR spectroscopy in transmission mode. A weighted minimization is used for the fitting.
Results. We present infrared spectra with corresponding band strengths of pure methane and binary mixtures with methane: CH4:H2O, CH4:CO2, CH4:CH3OH, CH4:NH3 at 6.7 K and 10 K showing a 20% increase in mixtures compared to commonly used 10 K band strength value of pure methane. We also test the usability of the spectra on open JWST data by probing the spatial distribution of methane in B335. We also present additional experiments concerning the phase transition of methane between phase II* and phase II.
Conclusions. Our results reveal distinct spectral features for methane in non-H2O environments, enabling more accurate interpretation of JWST observations. The dataset of spectra, publicly available on Zenodo, can be used for fitting JWST data.
Key Words.:
Astrochemistry, Methods: laboratory: solid state, Stars: protostars, ISM: abundances, ISM: molecules, Infrared: ISM1 Introduction
Solid methane is the most abundant and one of the simplest hydrocarbon molecules found in prestellar objects and protostars (Öberg et al., 2008). Both theoretical (Hassel et al., 2008; Harada and Herbst, 2008; Carder et al., 2021) and experimental (Baratta et al., 2002; Hodyss et al., 2009; Abplanalp et al., 2018; Zhu et al., 2018) studies consider methane a precursor for more complex carbonaceous molecules. Also, it can possibly be considered as a biosignature for exoplanetary studies (Krissansen-Totton et al., 2018; Thompson et al., 2022; Glass and Hörst, 2025). Methane is a part of the atmosphere of planets, e.g. Jupiter and Mars (Gautier et al., 1982; Krasnopolsky et al., 2004) and exoplanets, e.g. HD 209458b and HR 8799 b (Swain et al., 2009; Barman et al., 2015).
The first discovery of interstellar solid methane occurred by the ground-based infrared echelle spectrograph (IRSHELL) (Lacy et al., 1991). Methane is usually identified on infrared (IR) spectra using 7.7 m (1300 cm-1) deformation mode absorbance band. Boogert et al. (1996) performed the first fit of methane absorption band at 7.7 m with CH4:H2O ice spectrum and proposed the CH4:CH3OH and CH4:NH3 mixtures as alternatives. Based on shape and position of absorbance band it was concluded that methane is contained within the polar layer of ice mantles on interstellar dust (Boogert et al., 1998). Methane was identified in a number of objects in Infrared Space Observatory (ISO) data review (Gibb et al., 2004), where spectra of pure methane and methane in mixture with H2O were used. Later, methane abundance was constrained using Spitzer data, where methane was fitted with CH4:H2O mixtures (Öberg et al., 2008). Currently, research on methane continues with the recent James Webb Space Telescope (JWST) data. The deformation mode of solid methane at 7.7 m is described by methane-bearing H2O ices: at 15 K (H2O:CH4 = 10:1111Mixtures used in fitting procedures in the cited work are shown with this notation, Rocha et al., 2024) in IRAS 2A and IRAS 23385+6053, at 13 K (H2O:CH4 = 100:6, Chen et al., 2024) in B1-c, at 12 K (H2O:CH4 = 10:1, Rocha et al., 2025) in Ced 110 IRS4, H2O and CO2 environment at 12 K222Note that the spectrum that is referred to in fitting procedure is not listed in the original paper (Rocha et al., 2017) (H2O:CO2:CH4 = 10:1:1, McClure et al., 2023) in NIR38 and J110621, using a complex mixture at 10 K (H2O:CH3OH:CO2:CH4 = 0.6:0.7:1:0.1, Yang et al., 2022) in IRAS 15398–3359 and H2O:CH4 = 10:1 at 8 K and CO2:CH4 = 5:1 at 27 K (Nakibov et al., 2025) in IRAS 23385+6053. According to these findings on JWST data the methane abundance in cold clouds and star-forming regions is 1.6–3.7% with respect to solid H2O.
There are three phases of solid methane under vacuum conditions: phase I is stable above 20.4 K, phase II — below 20.4 K (e.g. Gerakines and Hudson, 2015b), and the third phase, stable if methane is deposited below 10 K (Colwell et al., 1962; Krupskii et al., 1973; Emtiaz et al., 2020). Following Emtiaz et al. (2020) we refer to this third phase as crystalline phase II*. However, the literature spectra of methane typically used for interpretation of observational data are obtained at a temperature of 10 K and above, and the laboratory data on ice analogues are present as water-based mixtures for phases I and II (Hudson and Moore, 1999; Ehrenfreund et al., 1999; Rocha et al., 2017; Mifsud et al., 2023). Binary mixtures of methane with H2O, CO, CH3OH, NH3, CO2, N2, O2 for the first time were described in Boogert et al. (1997) at 10 K yet the spectral data needed for fitting is unavailable. There are also reflection-mode studies (e.g. Kaiser et al., 2014; Maity et al., 2014; Förstel et al., 2017; Bergantini et al., 2018) that characterize methane photochemistry using relevant laboratory ices, however, their spectra are not directly comparable to astronomical observations of ices in transmission.
In this paper we present laboratory spectra of binary methane-containing ices in transmission regime, that are currently missing in the literature. These mixtures correspond to the environment of the polar layer of ice on the dust grains and aimed for interpretation of JWST data. The ices have astrochemically relevant component ratios: CH4:H2O = 1:10, CH4:CO2 = 1:5 and 1:15, CH4:CH3OH = 1:3, CH4:NH3 = 1:3. The set of mixtures is obtained at a substrate temperature of 10 K. However, in cold prestellar cores temperature may reach as low as 6 K, which is confirmed by, e.g. observations of gas and dust in centres of isolated cold dense cores L183 and L1544 (Pagani et al., 2004, 2007; Crapsi et al., 2007; Pagani et al., 2015). Also, physical modeling of L1544 resulted in the dust temperature in the core center equal to 6.5 K (Keto and Caselli, 2010). The presence of dust colder than 10 K is also confirmed by observations obtained with SCUBA at 850 m (Evans et al., 2001; Pattle et al., 2015). Therefore the set was extended with ices grown at 6.7 K with same composition. The extended set explores the difference between the ices grown at 6.7 K and 10 K substrate temperatures, which correspond to crystalline II* and crystalline II phases of pure methane. The transmission spectra of methane in binary mixtures grown at 6.7 K and CH4:CH3OH, CH4:NH3 at 10 K are presented here for the first time. The band strengths of methane in various environment are provided as well.
In our previous paper (Nakibov et al., 2025) we described gaseous and solid methane in IRAS 23385+6053 using some of the spectra presented here. In this paper we show the utility of our new infrared spectra by exploring the spatial distribution of methane in JWST spectra of the variable protostar B335-IRS (hereafter B335). B335 is an isolated nebula at a distance of 90–120 pc with a young class 0 protostar IRAS19347+0727 (age yr, Yıldız et al., 2015). It is associated with protostellar high-velocity Herbig-Haro (HH) jet (Reipurth et al., 1992; Gâlfalk and Olofsson, 2007). This source has a high level of accretion and two outflows (Bjerkeli et al., 2019) and had a luminosity increase by a factor of 5–7 (Evans et al., 2023). This source is a natural laboratory to study ice chemistry at the early stages of star formation. While the sublimation of complex ices and the rich gas-phase chemistry near the heated core was observed at millimeter wavelengths (Lee et al., 2025), we focus here on the solid methane features at 7.7 m located within the outflow cavity wall. This allows us to probe the composition of ices in the cold envelope of a young protostar.
2 Experimental methods
All experiments on growing analogues of interstellar ice were carried out on Ice Spectroscopy Experimental Aggregate (ISEAge) that utilizes ultra-high vacuum setup. A detailed description of our installation is provided in Ozhiganov et al. (2024). Briefly, the molecules are deposited on the Ge substrate from two sides via ‘background deposition’. The 6 mm diameter circle area (S=0.28 cm2) of the window is available for the deposition of the ice. The substrate is mounted on a cryogenic finger of a closed cycle He cryostat, and the whole chamber is continuously evacuated to the base pressure of × mbar. The operating temperatures are within the 6.7–305 K range with an error 0.1 K. The setup is equipped with a Stanford Research Systems RGA200 quadrupole mass spectrometer and Thermo Scientific Nicolet iS50 Fourier Transform Infrared (FTIR) spectrometer operating within a 4000–630 cm-1 (2.5–15.9 m) with 1 cm-1 resolution. The points in the collected spectra are spaced at 0.12 cm-1 intervals due to the Zero Filling performed on the interferometer level. The infrared spectra are collected in transmission regime. The FTIR spectrometer and optical path outside the main chamber are flushed by pure N2. Gaseous species are introduced into the main chamber via two independent leak valves.
We followed the same experimental procedure for growing of all ices. The component ratio was controlled during the deposition by means of the quadrupole mass spectrometer (QMS). In all experiments both components of a binary mixture were deposited simultaneously. The first component introduced into the main chamber was the pure methane and the second component was a matrix molecule introduced shortly after the desired ion current of methane was reached. The mixture deposition time started when ion currents for both components reached the desired values. This approach allows to correct the influence of ion interference and overlapping ion currents, and maintain the same ion current for methane across all the experiments. The ion current values were taken from the calibration curves obtained for our setup via the procedure similar to Slavicinska et al. (2023). Briefly, the rate at which gaseous molecules are adsorbed on the substrate depends nonlinearly on the ion current and is unique to each molecule. By growing pure ices at different ion currents we can obtain this dependency via experiment allowing for a more precise control over the components ratio of mixed ices. The rates for the calibration curves are determined by measuring the absorbance on the IR spectrum output using the literature band strength values. Since the depositions were conducted on the Ge substrate cooled to 10 K and 6.7 K we used the band strength for methane in phase II (corrected for density) from Bouilloud et al. (2015) and for methane in phase II* from Gerakines and Hudson (2015b), correspondingly.
For both sets of experiments the methane deposition rate was fixed at cm-2 s-1 and the second component deposition rate varied based on the desired components ratio. Deposition continued for 120 minutes. Thus, we get a column density of cm-2 for CH4. This approach enables the comparison of methane deformation (1300 cm-1) and stretching (3000 cm-1) modes in various molecular environment. The IR spectra were recorded every 45 seconds (an average of 32 scans). For the analysis in each experiment the last recorded spectrum of deposition was used to avoid the secondary adsorption effect after closing the leak valve. This effect is present when working with volatile species such as CH4, CO or N2 both in static conditions and during the warm up, see Appendix A for details. For calculations of peak positions and areas spectra are baseline-corrected for overall slope and local thin-film interference artifacts. Repeated measurements (50 iterations) with randomized local baseline adjustments are analyzed to minimize noise effects. The final values are derived from the averaged results.
The compounds used are as follows: CH4 (99.999 %, Ugra-PGS), CO2 (99.9999 %, Ugra-PGS), NH3 (99.9 %, UralKrioGaz), deionized H2O, CH3OH (99.8 %, Vekton). Gaseous CO2 and CH4 are introduced directly into the dosing lines from the commercially acquired gas bottles. H2O, NH3 and CH3OH were purified by a freeze-pump-thaw cycle three times before each experiment. Before each deposition every compound is checked for purity by means of mass-spectrometry.
3 Results
3.1 Binary mixtures with CH4
In this section we present a series of experiments on obtaining the binary ices of methane with molecules that constitute the polar layer of the interstellar ice mantles. We obtained binary mixtures at a temperature of 10 K, with methane molecules embedded in a matrix of H2O, CO2, CH3OH or NH3. The component ratio was chosen to qualitatively reflect current astrochemical view on species abundances in interstellar ices in cold dense molecular clouds (Boogert et al., 2015; Goto et al., 2021; McClure et al., 2023; Jiménez-Serra et al., 2025). Additionally, the comparative set was obtained at the minimum stable temperature of the ISEAge setup at 6.7 K. Methane absorption bands in mixtures at 6.7 K spectra differ both from pure methane and 10 K mixtures features. We present the spectra of mixtures for both temperatures in Fig. 1. Additional experiments concerning pure methane and its transition between phases II* and II are presented in Appendix A.
The methane absorption bands change shape and position in mixed ices, the components ratio and peak position of 3000 cm-1 stretching and 1300 cm-1 deformation mode are presented in Table 1. The band strengths for the deformation mode of methane in various mixtures are also included. Methane deformation mode in CH4:CO2 = 1:5 mixture stands out because of its structure. It displays an unresolved double peak both in 6.7 K and 10 K experiments, which becomes resolved on heating after 30–40 K (Boogert et al., 1997; Nakibov et al., 2025). Moreover, this mixture displays the most distinction between the two temperatures in both vibrational modes. This is caused by pure CO2 having an amorphous and crystalline form as shown in Gerakines and Hudson (2015a). For our fixed column density of 2.121017 cm-2 pure CO2 fully crystallizes at 10 K and only partially at 6.7 K. An addition of methane, however, prevents CO2 from crystallizing at 6.7 K. To further investigate this we obtained more diluted mixtures: CH4:CO2 = 1:15 at 6.7 K and 10 K, the differences persisted despite the higher CO2 concentration. We report, therefore, the peak shape of methane differs significantly in crystalline and amorphous CO2. We also note that the double peak gets resolved at 10 K in CH4:CO2 = 1:15 mixture, which is in line with Hudgins et al. (1993), where the double structure is observed in CH4:CO2 = 1:20 mixture at 10 K.
Methane mixtures with CH3OH and NH3 show similar features in 1300 cm-1 region and have close band strengths values. The features are redshifted by 2 cm-1 from the H2O:CH4 feature. This opens a possibility of singling out methane not related to water ice in observational spectra, an idea first introduced by Boogert et al. (1996). This is important, as normally methane is assumed to be formed simultaneously with water and ammonia ice via hydrogenation of carbon atoms at the onset of the formation of icy grain mantles (Qasim et al., 2020; Lamberts et al., 2022; Ferrero et al., 2023). However, methane can also be formed as a product of photoprocessing or ion-bombardment of methanol-rich ices (Islam et al., 2014) at later stages of ice evolution or via electron-induced radiolysis (Sullivan et al., 2016). The presented data allows potentially to explore the relative importance of both scenarios of methane formation.
| Mixture | Component ratio | Temperature (K) | Peak position (cm-1) | Peak position (cm-1) | Band Strength (cm) |
| CH4 | - | 8 | 3010.0 | 1296.9 | 1.04×10-17aaaaThe band strength is taken from Gerakines and Hudson (2015b); |
| CH4 | - | 10 | 3008.8 | 1301.3 | 8.4×10-18bbbbThe band strength is taken from Bouilloud et al. (2015). |
| 3020.5 | 1297.0 | ||||
| CH4:H2O | 1 : 10 | 6.7 | 3008.9 | 1302.7 | 8.8×10-18 |
| CH4:H2O | 1 : 10 | 10 | 3008.6 | 1302.5 | 1.0×10-17 |
| CH4:CO2 | 1 : 5 | 6.7 | 3010.6 | 1305.6 | 9.3×10-18 |
| CH4:CO2 | 1 : 5 | 10 | 3010.1 | 1301.7 | 1.2×10-17 |
| CH4:CO2 | 1 : 15 | 6.7 | 3011.7 | 1307.0 | 1.0×10-17 |
| CH4:CO2 | 1 : 15 | 10 | 3009.8 | 1308.6 | 1.1×10-17 |
| 3039.9 | 1300.1 | ||||
| CH4:CH3OH | 1 : 3 | 6.7 | 3005.6 | 1300.8 | 1.0×10-17 |
| CH4:CH3OH | 1 : 3 | 10 | 3005.2 | 1300.3 | 1.1×10-17 |
| CH4:NH3 | 1 : 3 | 6.7 | 3005.6 | 1300.2 | 1.0×10-17 |
| CH4:NH3 | 1 : 3 | 10 | 3007.0 | 1299.7 | 1.2×10-17 |
3.2 B335 methane survey
To test the usability of the obtained spectra we fitted the 7.7 m methane absorption band of the Class 0 protostar B335. JWST Mid-IR Instrument (MIRI) Medium Resolution Spectrograph (MRS) observations toward B335 are provided within the Investigating Protostellar Accretion proposal (PID 1802, P.I. T. Megeath; Megeath et al., 2021). Level 3 observation data were taken from the MAST database (MAST DOI: 10.17909/yr06-tf19), we further scaled the CH2-SHORT channel data to match the CH1-LONG signal level. MIRI MRS observations included three levels of JWST calibration pipeline with version 1.19.1 and Calibration References Data System (CRDS) context jwst_1413.pmap (CRDS_VER = ’12.1.11’) describing in Greenfield and Miller (2016); Bushouse et al. (2024); van Gelder et al. (2024). Unfortunately, the overlap region coincides with the OCN- absorption band and we excluded this specie from the further analysis.
B335 is viewed edge-on and has two prominent outflows (e.g. Bjerkeli et al., 2019), the west of which displaying strong continuum emission around 7.7 m. Based on the intensity map at 8 m (Fig. 2) we picked multiple apertures that cover bright pixels. This approach maximizes the signal-to-noise ratio, enhancing the analysis of the finer spectral details. To date, this is the first study that explores the spatial distribution on MIRI-MRS JWST data for a molecule with relatively low abundance relative to H2O. We note that a spatial distribution study was carried out recently for solid H2O, CO and CO2 on the NIRCam data (Smith et al., 2025). The spectra were extracted from the cylindrical apertures: seven 1.53′′ apertures were used to assess the spatial distribution of methane and a 4.59′′ aperture was used to ensure consistency between fits in smaller apertures. The center positions for the chosen apertures are listed in Table 2.
All spectra were processed in the same way. After baseline subtraction a linear combination of methane laboratory mixtures was fitted to the observational spectrum. The linear scale coefficients were obtained by minimizing , given as:
where and are observational and modeled optical depths, respectively, and is the point-wise standard deviation of the observational spectrum. The standard deviation was obtained from the residuals between the original optical depth spectrum and a smoothed curve obtained with locally weighted scatterplot smoothing (LOWESS, Cleveland, 1979) with a 24-point window. Column densities of H2O and CO2 were estimated based on 6 m and 15.2 m bands, respectively.
The fitted spectra are shown in the Fig. 3. The - observational spectra correspond to the - apertures from Fig. 2. Fitting results show that about 70% and 30% of methane budget in B335 are best described by methane in H2O and methane in CO2 environments, respectively (see Table 2). A minor part of methane is found in CH3OH environment, however, the detection at this level may be sensitive to e.g. baseline selection. Spectrum is excluded from the analysis due to low signal-to-noise ratio. The 1.53′′ apertures display variations in ice composition with CH4 in CO2 to CH4 in H2O relative abundances decreasing towards the protostar. There is also an overall increase in column density towards the core, likely explained by an increase of density towards the center.
The presence of CH4 in CO2 matrix provides some insights on ice formation at early stages of the protostar. An interesting correlation is observed between CH4 and CO2. The Spearman’s and p-values (Spearman, 1904) for CO2 and CH4 column densities (, =0.007) and abundances (, =0.003) indicate strong, statistically significant, positive correlation between these species. The correlations between CH4 and CO2 are known from as early as Spitzer methane survey (Öberg et al., 2008). Modeling presented in Garrod and Pauly (2011) shows that under certain conditions methane and CO2 can form concurrently, occupying the innermost layers of ice mantles. Ferrero et al. (2024) explored the reactivity of 3C-OH2 complex with various species, revealing barrierless reactions with atomic O and H. These reactions lead to production of CH2 and CO, which are the direct precursors to CH4 and CO2, respectively. While parallel formation of methane and CO2 is not largely explored, these results can be interpreted as evidence towards the concurrence for these processes.
| Aperture | Coordinates | (CH4) | CH4:H2O | CH4:CO2 | CH4:CH3OH | (CO2) | (H2O) | (CH4)/(H2O) |
| [1017 cm-2] | [%] | [%] | [%] | [1018 cm-2] | [1019 cm-2] | [%] | ||
| total | 19:37:01.0868 | 5.0 | 64 | 32 | 4 | 2.8 | 1.2 | 4.3 |
| +07:34:08.9309 | ||||||||
| a | 19:37:01.1384 | 3.8 | 65 | 35 | - | 2.5 | 1.0 | 3.8 |
| +07:34:10.3022 | ||||||||
| b | 19:37:01.0348 | 5.6 | 70 | 30 | - | 3.1 | 1.2 | 4.8 |
| +07:34:10.3022 | ||||||||
| c | 19:37:01.1892 | 2.8 | 37 | 53 | 10 | 2.4 | 1.1 | 2.5 |
| +07:34:08.9298 | ||||||||
| d | 19:37:01.0868 | 4.1 | 52 | 34 | 14 | 2.2 | 1.2 | 3.4 |
| +07:34:08.9309 | ||||||||
| e | 19:37:00.9841 | 6.4 | 75 | 23 | 2 | 3.9 | 1.9 | 3.4 |
| +07:34:08.9298 | ||||||||
| f | 19:37:01.1375 | 4.1 | 74 | 26 | - | 2.4 | 1.3 | 3.1 |
| +07:34:07.5764 | ||||||||
| g | 19:37:01.0350 | 5.4 | 70 | 30 | - | 3.0 | 1.8 | 3.1 |
| +07:34:07.5764 |
4 Astrochemical implications and Conclusions
In this paper we present spectra of pure methane and binary mixtures with methane for 6.7 K and 10 K deposition temperatures.
The following mixtures are presented in this paper: CH4:H2O = 1:10, CH4:CO2 = 1:5, CH4:CO2 = 1:15, CH4:CH3OH = 1:3 and CH4:NH3 = 1:3. All spectra of methane in astrochemically relevant mixture ratios are presented for the first time at 6.7 K. The new band strength values and spectra obtained at temperature below 9 K reveal the following astrochemical implications:
1. Currently, the band strength of pure methane at 10 K is used majorly for column density estimations in observational data (e.g. McClure et al., 2023; Rocha et al., 2024, 2025). However, solid interstellar methane is found in mixtures with other icy species rather than in pure form. We also note that interstellar ices differ from laboratory analogues both in how they are produced and how they are processed. Thus, while the band strengths in Table 1 provide a better basis for interpreting observations, they are still approximate. Nevertheless, our new data invites a discussion on recent literature estimates. Since the methane band strengths in mixtures are 20 % higher on average than the band strength of pure methane at 10 K, the column densities of interstellar methane may be overestimated by this value in e.g. McClure et al. (2023). Interestingly, this partially explains the underestimation of solid methane abundance by astrochemical models toward background stars NIR38 and J110621 reported in Jiménez-Serra et al. (2025).
2. We present the the first publicly available digital spectral library of methane absorbance spectra in non-H2O environment, providing the possibility to identify non-H2O methane environments on newly obtained JWST spectra. Solid methane is typically considered to reside in polar water-rich environment due its early formation along with water ice (see e.g. Hasegawa et al., 1992; Clément et al., 2023; Borshcheva et al., 2025). This is further supported by successful fits of methane features in JWST spectra of interstellar ices with CH4:H2O mixtures (e.g. Chen et al., 2024; Rocha et al., 2024, 2025). However, there are possible evidences for its presence in mixtures with other molecules. For example, works by Boogert et al. (1996) and Nakibov et al. (2025) show that observed shapes of 7.7 m methane feature can be partially interpreted with non-H2O methane-containing mixtures. Additionally, icy methane mixed with apolar species such as N2 and CO is found on icy bodies in the outer Solar system (e.g., Emtiaz et al., 2022). The obtained spectra of ices of astrochemical ratios are published on Zenodo (DOI: 10.5281/zenodo.15350566). Comparison between spectra obtained at 6.7 K and 10 K reveal temperature-induced changes in shape and peak position, although we find it unlikely that these differences are prominent enough to be discerned within the observational data. However, it is known that central regions of prestellar cores are well-shielded from external heating radiation. Consequently, as shown both observationally (Pagani et al., 2004, 2007; Crapsi et al., 2007; Pagani et al., 2015) and theoretically (Keto and Field, 2005; Keto and Caselli, 2010, and references therein), dust temperature there may reach 6 K. According to a recent modeling by Borshcheva et al. (2025), approximately 80% of solid methane in the prestellar core L1544 resides on grains with temperature below 9 K. Thus, to accurately interpret future observations of ices on a line of sight toward L1544, laboratory spectra and band strengths of mixtures obtained both at 6.7 K and 10 K might be needed. The presented spectra can be directly compared to observational data, because the deformation mode of methane is unlikely to be affected by grain size, depending mostly on temperature and molecular environment (Boogert et al., 1997).
3. We present the first spatial survey performed on JWST MIRI-MRS data, exploring methane in B335. The results show that about 30% of solid methane is found in CO2 environment, rather than in H2O matrix. The consistency between findings in full and smaller apertures proves the usability of presented spectra for JWST data analysis. The observed ice composition gradient further highlights that JWST provides enough sensitivity for such studies, prompting further research.
5 Data availability
The laboratory infrared absorbance spectra in transmission regime in the 2.5–15.9 m (4000–630 cm-1) region for CH4 in different environments at 6.7 K and 10 K can be found on Zenodo at the following link: https://zenodo.org/records/15350566.
Acknowledgements.
The authors thank Prof. Dr. Paola Caselli for the discussion on the dust temperature in prestellar cores, and Ms. Katerina Borshcheva for providing the information on the distribution of methane ice in the model of L1544. We thank the anonymous reviewer for their comments that helped us to improve the manuscript. This research work is funded by the Russian Science Foundation via 23-12-00315 agreement.References
- Untangling the methane chemistry in interstellar and solar system ices toward ionizing radiation: a combined infrared and reflectron time-of-flight analysis. PCCP 20, pp. 5435–5468. External Links: Document, ADS entry Cited by: §1.
- Formation of a CO-CH 4 complex in thin solid films below 50 K. CPL 378 (1), pp. 178–184. External Links: Document, ADS entry Cited by: Appendix A.
- A comparison of ion irradiation and UV photolysis of CH4 and CH3OH. A&A 384, pp. 343–349. External Links: Document, ADS entry Cited by: §1.
- Simultaneous detection of water, methane, and carbon monoxide in the atmosphere of exoplanet hr 8799 b. ApJ 804 (1), pp. 61. External Links: Document Cited by: §1.
- A Mechanistical Study on the Formation of Dimethyl Ether (CH3OCH3) and Ethanol (CH3CH2OH) in Methanol-containing Ices and Implications for the Chemistry of Star-forming Regions. ApJ 852 (2), pp. 70. External Links: Document, ADS entry Cited by: §1.
- Kinematics around the B335 protostar down to au scales. A&A 631, pp. A64. External Links: Document, 1909.05178, ADS entry Cited by: §1, §3.2.
- The gas/solid methane abundance ratio toward deeply embedded protostars. A&A 336, pp. 352–358. External Links: Link Cited by: §1.
- Infrared observations and laboratory simulations of interstellar CH_4_ and SO_2_.. A&A 317, pp. 929–941. External Links: Link Cited by: Appendix A, §1, §3.1, §4.
- Solid methane toward deeply embedded protostars.. A&A 315, pp. L377–L380. External Links: ADS entry Cited by: §1, §3.1, §4.
- Observations of the icy universe.. ARA&A 53, pp. 541–581. External Links: Document, 1501.05317, ADS entry Cited by: §3.1.
- Formation of Complex Organic Molecules in Prestellar Cores: The Role of Nondiffusive Grain Chemistry. ApJ 990 (2), pp. 163. External Links: Document, 2507.00338, ADS entry Cited by: §4.
- Bibliographic review and new measurements of the infrared band strengths of pure molecules at 25 K: H2O, CO2, CO, CH4, NH3, CH3OH, HCOOH and H2CO. MNRAS 451 (2), pp. 2145–2160. External Links: Document, ADS entry Cited by: §2, footnote .
- JWST calibration pipeline External Links: Document, Link Cited by: §3.2.
- Modelling the insertion of O(1D) into methane on the surface of interstellar ice mantles. MNRAS 508 (1), pp. 1526–1532. External Links: Document, ADS entry Cited by: §1.
- JOYS+: The link between the ice and gas of complex organic molecules: Comparing JWST and ALMA data of two low-mass protostars. A&A 690, pp. A205. External Links: Document, 2407.20066, ADS entry Cited by: §1, §4.
- Astrochemical models of interstellar ices: History matters. A&A 675, pp. A165. External Links: Document, 2306.08346, ADS entry Cited by: §4.
- Robust locally weighted regression and smoothing scatterplots. JASA 74 (368), pp. 829–836. External Links: Document, Link, https://www.tandfonline.com/doi/pdf/10.1080/01621459.1979.10481038 Cited by: §3.2.
- A laboratory survey of the thermal desorption of astrophysically relevant molecules. MNRAS 354 (4), pp. 1133–1140. External Links: Document, ADS entry Cited by: Appendix A.
- Second Transition in Solid CH4. J. Chem. Phys. 36 (8), pp. 2223–2224. External Links: Document, ADS entry Cited by: Appendix A, Appendix A, §1.
- Observing the gas temperature drop in the high-density nucleus of L 1544. A&A 470 (1), pp. 221–230. External Links: Document, 0705.0471, ADS entry Cited by: §1, §4.
- Cosmic ray impact on astrophysical ices: laboratory studies on heavy ion irradiation of methane. A&A 531, pp. A160. External Links: Document, ADS entry Cited by: Appendix A.
- Laboratory studies of thermally processed H_2O-CH_3OH-CO_2 ice mixtures and their astrophysical implications. A&A 350, pp. 240–253. External Links: Link Cited by: §1.
- Infrared Spectroscopic Study of Solid Methane: Nuclear Spin Conversion of Stable and Metastable Phases. JPCA 124 (3), pp. 552–559. External Links: Document, ADS entry Cited by: Appendix A, Appendix A, §1.
- Infrared spectroscopic study of methane ice, pure and in mixtures with polar (h2o) and nonpolar (n2) molecules. JPCA 126 (12), pp. 1973–1979. External Links: Document, Link, https://doi.org/10.1021/acs.jpca.2c00287 Cited by: §4.
- Tracing the Mass during Low-Mass Star Formation. II. Modeling the Submillimeter Emission from Preprotostellar Cores. ApJ 557 (1), pp. 193–208. External Links: Document, astro-ph/0104238, ADS entry Cited by: §1.
- Models of Rotating Infall for the B335 Protostar. ApJ 943 (2), pp. 90. External Links: Document, 2212.03746, ADS entry Cited by: §1.
- Formation of Interstellar Complex Organic Molecules on Water-rich Ices Triggered by Atomic Carbon Freezing. ApJ 960 (1), pp. 22. External Links: Document, 2310.14831, ADS entry Cited by: §3.2.
- Where Does the Energy Go during the Interstellar NH3 Formation on Water Ice? A Computational Study. ApJ 944 (2), pp. 142. External Links: Document, 2212.14271, ADS entry Cited by: §3.1.
- Formation of Methylamine and Ethylamine in Extraterrestrial Ices and Their Role as Fundamental Building Blocks of Proteinogenic -amino Acids. ApJ 845 (1), pp. 83. External Links: Document, ADS entry Cited by: §1.
- Herbig-Haro flows in B335. A&A 475 (1), pp. 281–300. External Links: Document, 0709.4047, ADS entry Cited by: §1.
- On the Formation of CO2 and Other Interstellar Ices. ApJ 735 (1), pp. 15. External Links: Document, 1106.0540, ADS entry Cited by: §3.2.
- The C/H ratio in Jupiter from the Voyager infrared investigation. ApJ 257, pp. 901–912. External Links: Document, ADS entry Cited by: §1.
- Ultraviolet processing of interstellar ice analogs. I. Pure ices.. A&A 312, pp. 289–305. External Links: Link Cited by: Appendix A.
- First Infrared Band Strengths for Amorphous CO2, an Overlooked Component of Interstellar Ices. ApJ 808 (2), pp. L40. External Links: Document, ADS entry Cited by: §3.1.
- Infrared Spectra and Optical Constants of Elusive Amorphous Methane. ApJ 805 (2), pp. L20. External Links: Document, ADS entry Cited by: Appendix A, Appendix A, Appendix A, Appendix A, Appendix A, §1, §2, footnote .
- A Modified Algorithm and Open-source Computational Package for the Determination of Infrared Optical Constants Relevant to Astrophysics. ApJ 901 (1), pp. 52. External Links: Document, ADS entry Cited by: Appendix A.
- Interstellar Ice: The Infrared Space Observatory Legacy. ApJS 151 (1), pp. 35–73. External Links: Document, ADS entry Cited by: §1.
- The Once and Future Gas: Methane’s Multifunctional Roles in Earth’s Evolution and Potential as a Biosignature. AREPS 53 (1), pp. 283–304. External Links: Document, ADS entry Cited by: §1.
- Water and methanol ice in L 1544. A&A 651, pp. A53. External Links: Document, 2012.10883, ADS entry Cited by: §3.1.
- The Calibration Reference Data System. A&C 16, pp. 41–53. External Links: Document, ADS entry Cited by: §3.2.
- Modeling Carbon Chain Anions in L1527. ApJ 685 (1), pp. 272–280. External Links: Document, 0808.1249, ADS entry Cited by: §1.
- Models of Gas-Grain Chemistry in Dense Interstellar Clouds with Complex Organic Molecules. ApJS 82, pp. 167. External Links: Document, ADS entry Cited by: §4.
- Modeling the Lukewarm Corino Phase: Is L1527 Unique?. ApJ 681 (2), pp. 1385–1395. External Links: Document, 0803.1805, ADS entry Cited by: §1.
- Photochemistry of methane water ices. Icarus 200 (1), pp. 338–342. External Links: Document, ADS entry Cited by: §1.
- Mid- and Far-Infrared Spectroscopy of Ices: Optical Constants and Integrated Absorbances. ApJS 86, pp. 713. External Links: Document, ADS entry Cited by: Appendix A, §3.1.
- Activation of weak IR fundamentals of two species of astrochemical interest in the Tdpoint group - the importance of amorphous ices. PCCP 17 (19), pp. 12545–12552. External Links: Document, ADS entry Cited by: Appendix A, Appendix A.
- Laboratory Studies of the Formation of Methanol and Other Organic Molecules by Water+Carbon Monoxide Radiolysis: Relevance to Comets, Icy Satellites, and Interstellar Ices. Icar 140 (2), pp. 451–461. External Links: Document, ADS entry Cited by: §1.
- Simultaneous UV- and ion processing of astrophysically relevant ices. The case of CH3OH:N2 solid mixtures. A&A 561, pp. A73. External Links: Document, ADS entry Cited by: §3.1.
- Modelling methanol and hydride formation in the JWST Ice Age era. A&A 695, pp. A247. External Links: Document, 2502.10123, ADS entry Cited by: §3.1, §4.
- Infrared and reflectron time-of-flight mass spectroscopic analysis of methane (CH4)-carbon monoxide (CO) ices exposed to ionization radiation - toward the formation of carbonyl-bearing molecules in extraterrestrial ices. PCCP 16 (8), pp. 3399. External Links: Document, ADS entry Cited by: §1.
- Dynamics and depletion in thermally supercritical starless cores. MNRAS 402 (3), pp. 1625–1634. External Links: Document, 0908.2400, ADS entry Cited by: §1, §4.
- Dark Cloud Cores and Gravitational Decoupling from Turbulent Flows. ApJ 635 (2), pp. 1151–1165. External Links: Document, astro-ph/0508527, ADS entry Cited by: §4.
- Theory of phase transitions in solid methanes. XI. Infrared and Raman spectra of the 3 and 4 modes in phase II of solid CH4. J. Chem. Phys. 66 (12), pp. 5568–5577. External Links: Document, ADS entry Cited by: Appendix A, Appendix A.
- Detection of methane in the martian atmosphere: evidence for life?. Icarus 172 (2), pp. 537–547. External Links: Document, ADS entry Cited by: §1.
- Detectability of Biosignatures in Anoxic Atmospheres with the James Webb Space Telescope: A TRAPPIST-1e Case Study. AJ 156 (3), pp. 114. External Links: Document, 1808.08377, ADS entry Cited by: §1.
- Experimental observation of second phase transition in solid ch4. JETPL 17, pp. 605. External Links: Document, Link Cited by: Appendix A, Appendix A, §1.
- Discovery of Interstellar Methane: Observations of Gaseous and Solid CH 4 Absorption toward Young Stars in Molecular Clouds. ApJ 376, pp. 556. External Links: Document, ADS entry Cited by: §1.
- Methane Formation in Cold Regions from Carbon Atoms and Molecular Hydrogen. ApJ 928 (1), pp. 48. External Links: Document, 2110.15881, ADS entry Cited by: §3.1.
- A Natural Laboratory for Astrochemistry: The Variable Protostar B335. ApJ 978 (1), pp. L3. External Links: Document, 2410.05904, ADS entry Cited by: §1.
- Formation of Ketene (H2CCO) in Interstellar Analogous Methane (CH4)-Carbon Monoxide (CO) Ices: A Combined Ftir and Reflectron Time-of-Flight Mass Spectroscopic Study. ApJ 789 (1), pp. 36. External Links: Document, ADS entry Cited by: §1.
- An Ice Age JWST inventory of dense molecular cloud ices. NatAs 7, pp. 431–443. External Links: Document, 2301.09140, ADS entry Cited by: §1, §3.1, §4.
- Investigating Protostellar Accretion Across the Mass Spectrum. Note: JWST Proposal. Cycle 1, ID. #1802 External Links: ADS entry Cited by: §3.2.
- Proton and Electron Irradiations of CH4:H2O Mixed Ices. Atoms 11 (2), pp. 19. External Links: Document, 2301.09102, ADS entry Cited by: §1.
- X-ray Diffraction Study on Simple Molecular Glasses Created by Low-Temperature Vapor Deposition. JPSJ 85 (12), pp. 124602. External Links: Document, ADS entry Cited by: Appendix A.
- Intermolecular correlations of liquid and glassy CS2 studied by synchrotron radiation x-ray diffraction. J. Chem. Phys. 156 (3), pp. 034503. External Links: Document, ADS entry Cited by: Appendix A.
- Profile of CH _4 IR bands in ice mixtures. A&A 333, pp. 1025–1033. External Links: Link Cited by: Appendix A.
- Solid and Gaseous Methane in IRAS 23385+6053 as Seen with Open JWST Data. ApJ 978 (2), pp. L46. External Links: Document, 2412.17028, ADS entry Cited by: §1, §1, §3.1, §4.
- The c2d Spitzer Spectroscopic Survey of Ices around Low-Mass Young Stellar Objects. III. CH4. ApJ 678 (2), pp. 1032–1041. External Links: Document, 0801.1223, ADS entry Cited by: §1, §1, §3.2.
- Infrared Spectra of Solid HCN Embedded in Various Molecular Environments for Comparison with the Data Obtained with JWST. ApJ 972 (1), pp. L10. External Links: Document, 2408.08166, ADS entry Cited by: §2.
- Depletion and low gas temperature in the L183 (=L134N) prestellar core: the N2H+̂-N2D+ tool. A&A 467 (1), pp. 179–186. External Links: Document, astro-ph/0701823, ADS entry Cited by: §1, §4.
- L183 (L134N) Revisited. II. The dust content. A&A 417, pp. 605–613. External Links: Document, ADS entry Cited by: §1, §4.
- Can we trace very cold dust from its emission alone?. A&A 574, pp. L5. External Links: Document, 1501.00861, ADS entry Cited by: §1, §4.
- The JCMT Gould Belt Survey: first results from the SCUBA-2 observations of the Ophiuchus molecular cloud and a virial analysis of its prestellar core population. MNRAS 450 (1), pp. 1094–1122. External Links: Document, 1502.05858, ADS entry Cited by: §1.
- An experimental study of the surface formation of methane in interstellar molecular clouds. NatAs 4, pp. 781–785. External Links: Document, 2004.02506, ADS entry Cited by: §3.1.
- Star formation in Bok globules and low-mass clouds. IV. Herbig-Haro objects in B335.. A&A 256, pp. 225. External Links: ADS entry Cited by: §1.
- Ice inventory towards the protostar Ced 110 IRS4 observed with the James Webb Space Telescope: Results from the Early Release Science Ice Age program. A&A 693, pp. A288. External Links: Document, 2411.19651, ADS entry Cited by: §1, §4, §4.
- Infrared complex refractive index of astrophysical ices exposed to cosmic rays simulated in the laboratory. MNRAS 464 (1), pp. 754–767. External Links: Document, ADS entry Cited by: §1, footnote 2.
- JWST Observations of Young protoStars (JOYS+): Detecting icy complex organic molecules and ions. I. CH4, SO2, HCOO-, OCN-, H2CO, HCOOH, CH3CH2OH, CH3CHO, CH3OCHO, and CH3COOH. A&A 683, pp. A124. External Links: Document, 2312.06834, ADS entry Cited by: §1, §4, §4.
- Studying pure ch4 and the interaction of ch4 and h2o in interstellar ice analogues with on-resonance infrared irradiation. JPCA 129 (30), pp. 6883–6895. Note: PMID: 40680293 External Links: Document, Link, https://doi.org/10.1021/acs.jpca.5c03186 Cited by: Appendix A.
- The hunt for formamide in interstellar ices. A toolkit of laboratory infrared spectra in astronomically relevant ice mixtures and comparisons to ISO, Spitzer, and JWST observations. A&A 677, pp. A13. External Links: Document, 2307.04790, ADS entry Cited by: §2.
- Cospatial ice mapping of H2O with CO2 and CO across a molecular cloud with JWST/NIRCam. NatAs 9, pp. 883–894. External Links: Document, ADS entry Cited by: §3.2.
- The proof and measurement of association between two things. The American Journal of Psychology 15 (1), pp. 72–101. External Links: ISSN 00029556, Link, Document Cited by: §3.2.
- Low-energy (¡20 eV) and high-energy (1000 eV) electron-induced methanol radiolysis of astrochemical interest. MNRAS 460 (1), pp. 664–672. External Links: Document, ADS entry Cited by: §3.1.
- Water, Methane, and Carbon Dioxide Present in the Dayside Spectrum of the Exoplanet HD 209458b. ApJ 704 (2), pp. 1616–1621. External Links: Document, 0908.4010, ADS entry Cited by: §1.
- The case and context for atmospheric methane as an exoplanet biosignature. PNAS 119 (14), pp. e2117933119. External Links: Document, Link, https://www.pnas.org/doi/pdf/10.1073/pnas.2117933119 Cited by: §1.
- JOYS+: Mid-infrared detection of gas-phase SO2 emission in a low-mass protostar. The case of NGC 1333 IRAS 2A: Hot core or accretion shock?. A&A 682, pp. A78. External Links: Document, 2311.17161, ADS entry Cited by: §3.2.
- CORINOS. I. JWST/MIRI Spectroscopy and Imaging of a Class 0 Protostar IRAS 15398–3359. ApJ 941 (1), pp. L13. External Links: Document, 2208.10673, ADS entry Cited by: §1.
- APEX-CHAMP+ high-J CO observations of low-mass young stellar objects. IV. Mechanical and radiative feedback. A&A 576, pp. A109. External Links: Document, 1501.03259, ADS entry Cited by: §1.
- Formation and High-order Carboxylic Acids (RCOOH) in Interstellar Analogous Ices of Carbon Dioxide (CO2) and Methane(CH4). ApJS 234 (1), pp. 15. External Links: Document, ADS entry Cited by: §1.
Appendix A pure CH4
There are three phases of solid methane under vacuum conditions. The crystalline phase II is stable below 20.4 K and has the fcc lattice with eight sublattices: six of which have site symmetry D2d and two — site symmetry Oh (Kobashi et al. 1977). It displays a double peak structure of both (3009 cm-1 and 3021 cm-1) and (1301 cm-1 and 1298 cm-1) features. This phase reversibly transitions into phase I above the first transition temperature of 20.4 K and has fcc lattice with a single freely rotating molecule. The and features are positioned at 3010 cm-1 and 1299 cm-1, respectively. Phase II* is observed if methane ice is obtained by vapor deposition below 10 K. It displays a degree of orientational disorder between phase I and phase II (Emtiaz et al. 2020). In this phase the absorption bands of methane are positioned near 3010 cm-1 and 1297 cm-1 (Gerakines and Hudson 2015b). Phase II* cannot be obtained by cooling methane ice from the triple point or from phase II ice. On warm-up phase II* ice undergoes irreversible transition to phase II. The second transition temperature depends on experimental conditions due to metastability and is observed below 10 K (Colwell et al. 1962; Krupskii et al. 1973; Emtiaz et al. 2020).
Despite the extensive research carried out on methane solid phases in infrared at low temperatures in recent decades (Kobashi et al. 1977; Hudgins et al. 1993; Boogert et al. 1997; Mulas et al. 1998; Gerakines and Hudson 2015b; Hudson et al. 2015; Emtiaz et al. 2020; Gerakines and Hudson 2020), there is still an ambiguity between published studies. In several works (Hudgins et al. 1993; Gerakines et al. 1996; Mulas et al. 1998; Alsindi et al. 2003; de Barros et al. 2011) methane is claimed to be in phase II, but the presented spectral features differ significantly from the theoretically derived shape (Kobashi et al. 1977). Gerakines and Hudson (2015b) addressed this discrepancy by providing spectra of ‘amorphous methane’, suggesting that the literature spectra in the works above display partial amorphous-to-crystalline transition, which was caused by the high deposition rates and experimental constraints of the setups used.
Gerakines and Hudson (2015b) report that the majority of their spectra were obtained in the 8–12 K range, but the spectra are referred to as obtained ‘at 10 K’ for simplicity throughout the text. Thus, it is not clear at what temperature exactly were published experiments carried out, a point that also applies to Hudson et al. (2015). This ambiguity is significant, because the 8–12 K range includes published methane second transition temperatures (Colwell et al. 1962; Krupskii et al. 1973). We attempted to obtain the spectrum at their Fig. 1(f) at precisely 10 K, but failed even using a lower deposition rate than stated in their paper. Therefore, prior to obtaining the set of methane-bearing mixtures at 10 K we decided to investigate the transition between phase II* and phase II. Under the deposition rate and duration of experiments described in the main text methane retained phase II* structure when deposited at 6.7 K throughout the whole deposition time. In contrast, for deposition at 10 K, the double peak structure characteristic of phase II started to form once ice reached certain column density, see Figure 4. At a deposition rate of 5.9 cm-2 s-1 the transition started after 35 minutes, corresponding to a column density of 1.21016 cm-2. We also conducted depositions at rates twice as low and twice as high: 2.95 cm-2 s-1 and 1.18 cm-2 s-1. In all cases, on further deposition the double peak evolved through intermediate shapes, similar to those shown by Gerakines and Hudson (2015b), before fully converting to the phase II structure. The change in peak structure occurs at a critical value of 1.21016 cm-2. This equals approximately to 6 monolayers of pure methane ice on each side of the Ge substrate, given that single monolayer column density is 11015 cm-2.
From our experiments we conclude that phase II* is stable at 6.7 K. At 10 K phase II* is only observed in very thin methane films. In our experiments on Ge substrate, the maximum ice column density at which phase II* is still observed at 10 K is 1.21016 cm-2, independent of the deposition rate. Further ice growth at 10 K leads to the transition of phase II* to phase II through intermediate spectral shapes. While these intermediate spectra are similar to those mentioned in Gerakines and Hudson (2015b) we refrain from a more direct comparison. We expect that similar results can be obtained under minor differences in experimental conditions of other setups, which is expected given the environment-sensitive nature of metastable phase transition. A definitive answer would come from X-ray diffractometry study as performed in Mizuno et al. (2016, 2022), which showed that other molecular ices that are ‘conventionally amorphous’ for astrochemistry can exhibit positionally ordered but orientationally disordered structure (CO or N2), while other (CO2, CS2) are indeed amorphous.
We also note an instrumental effect that concerns both temperature programmed desorption (TPD) and static temperature experiments with volatile species. Here, we interpret the cases transparently reported by Collings et al. (2004) and Schrauwen et al. (2025). Methane temporarily adsorbs on warm surfaces (some cryostat parts, radiation shield, etc.) during the deposition and continuously releases even when the leak valves are closed, as in Schrauwen et al. (2025). The evacuation therefore is governed by the amount of temporarily stored methane rather than by turbomolecular pump’s efficiency, resulting in a ‘tail’ in QMS signal after deposition. During TPD warm surfaces release stored material early, polluting the QMS signal, as in Collings et al. (2004). Released gas can get readsorbed by the cold substrate. An example of that can be found in the cyclic experiment in Gerakines and Hudson (2015b) (see their Fig. 1, right column). We note that the integrated intensity increased after the first heating-holding-cooling cycle, indicating an increase in column density. Second cycle did not cause an increase as stored methane depleted on first heating. This increase did not affect the results of their paper.

### v2

@@
-Infrared spectra of methane-containing ice mixtures for JWST data analysis
+Infrared spectra of methane-containing ice mixtures for JWST data analyses
 Abstract
 Context.
-Solid methane (CH4) is an important molecule in interstellar and planetary environments, serving as a precursor to complex organic compounds and a potential biosignature in exoplanetary studies.
-Despite its significance, laboratory data on low-temperature phase of methane below 10 K remain limited.
+Solid methane (CH4) is an important molecule in interstellar and planetary environments, serving as a precursor to complex organic compounds, and it is a potential biosignature in exoplanetary studies.
+Despite its significance, laboratory data on the low-temperature phase of methane below 10 K remain limited.
 Aims.
-We aim to obtain spectra of methane in binary mixtures at 10 K and compare it to the spectra obtained at 6.7 K.
-These temperatures correspond to phases II and II* of pure methane and are representative of dark molecular clouds and protostars at early stages.
-We also aim to test the obtained data applicability to JWST data interpretation.
-Methods. Laboratory reference spectra were obtained on the ISEAge setup via FTIR spectroscopy in transmission mode. A weighted minimization is used for the fitting.
+We obtained spectra of methane in binary mixtures at 10 K and compared them to the spectra obtained at 6.7 K.
+These temperatures correspond to phases II and II* of pure methane and are representative of dark molecular clouds and protostars in early stages.
+We also tested whether the data we obtained can be applied to interpret JWST data.
+Methods. Laboratory reference spectra were obtained with the ISEAge setup via Fourier transform infrared spectroscopy in transmission mode. A weighted minimization was used for the fitting.
 Results.
-We present infrared spectra with corresponding band strengths of pure methane and binary mixtures with methane: CH4:H2O, CH4:CO2, CH4:CH3OH, CH4:NH3 at 6.7 K and 10 K showing a 20% increase in mixtures compared to commonly used 10 K band strength value of pure methane.
-We also test the usability of the spectra on open JWST data by probing the spatial distribution of methane in B335.
+We present infrared spectra with corresponding band strengths of pure methane and binary mixtures with methane: CH4:H2O, CH4:CO2, CH4:CH3OH, and CH4:NH3 at 6.7 K and 10 K.
+They show an increase of 20% in mixtures compared to the commonly used 10 K band strength value of pure methane.
+We also tested whether the spectra can be used on open JWST data by probing the spatial distribution of methane in B335.
 We also present additional experiments concerning the phase transition of methane between phase II* and phase II.
 Conclusions.
-Our results reveal distinct spectral features for methane in non-H2O environments, enabling more accurate interpretation of JWST observations.
-The dataset of spectra, publicly available on Zenodo, can be used for fitting JWST data.
+Our results reveal distinct spectral features for methane in non-H2O environments that enable a more accurate interpretation of JWST observations.
+The dataset of spectra is publicly available on Zenodo and can be used for fitting JWST data.
 Key Words.:
 Astrochemistry, Methods: laboratory: solid state, Stars: protostars, ISM: abundances, ISM: molecules, Infrared: ISM1 Introduction
-Solid methane is the most abundant and one of the simplest hydrocarbon molecules found in prestellar objects and protostars (Öberg et al., 2008).
-Both theoretical (Hassel et al., 2008; Harada and Herbst, 2008; Carder et al., 2021) and experimental (Baratta et al., 2002; Hodyss et al., 2009; Abplanalp et al., 2018; Zhu et al., 2018) studies consider methane a precursor for more complex carbonaceous molecules.
-Also, it can possibly be considered as a biosignature for exoplanetary studies (Krissansen-Totton et al., 2018; Thompson et al., 2022; Glass and Hörst, 2025).
-Methane is a part of the atmosphere of planets, e.g.
-Jupiter and Mars (Gautier et al., 1982; Krasnopolsky et al., 2004) and exoplanets, e.g.
-HD 209458b and HR 8799 b (Swain et al., 2009; Barman et al., 2015).
-The first discovery of interstellar solid methane occurred by the ground-based infrared echelle spectrograph (IRSHELL) (Lacy et al., 1991).
-Methane is usually identified on infrared (IR) spectra using 7.7 m (1300 cm-1) deformation mode absorbance band.
+Solid methane is the most abundant and one of the simplest hydrocarbon molecules found in pre-stellar objects and protostars (Öberg et al., 2008).
+Both theoretical (Hassel et al., 2008; Harada and Herbst, 2008; Carder et al., 2021) and experimental (Baratta et al., 2002; Hodyss et al., 2009; Abplanalp et al., 2018; Zhu et al., 2018) studies have considered methane a precursor for more complex carbonaceous molecules.
+It is also considered a biosignature for exoplanetary studies (Krissansen-Totton et al., 2018; Thompson et al., 2022; Glass and Hörst, 2025).
+Methane is a part of the atmosphere of planets, for instance Jupiter and Mars (Gautier et al., 1982; Krasnopolsky et al., 2004), and exoplanets, for example HD 209458b and HR 8799 b (Swain et al., 2009; Barman et al., 2015).
+The first discovery of interstellar solid methane was made with the ground-based infrared (IR) echelle spectrograph IRSHELL (Lacy et al.
+1991).
+Methane is usually identified on IR spectra using the 7.7 m (1300 cm-1) deformation mode absorbance band.
 Boogert et al.
-(1996) performed the first fit of methane absorption band at 7.7 m with CH4:H2O ice spectrum and proposed the CH4:CH3OH and CH4:NH3 mixtures as alternatives.
-Based on shape and position of absorbance band it was concluded that methane is contained within the polar layer of ice mantles on interstellar dust (Boogert et al., 1998).
-Methane was identified in a number of objects in Infrared Space Observatory (ISO) data review (Gibb et al., 2004), where spectra of pure methane and methane in mixture with H2O were used.
-Later, methane abundance was constrained using Spitzer data, where methane was fitted with CH4:H2O mixtures (Öberg et al., 2008).
+(1996) performed the first fit of the methane absorption band at 7.7 m with the CH4:H2O ice spectrum and proposed the CH4:CH3OH and CH4:NH3 mixtures as alternatives.
+Based on the shape and position of the absorbance band, it was concluded that methane is contained within the polar layer of ice mantles on interstellar dust (Boogert et al., 1998).
+Methane was identified in a number of objects in a review of Infrared Space Observatory data (Gibb et al., 2004), where spectra of pure methane and methane mixed with H2O were used.
+Later, the methane abundance was constrained using Spitzer data, where methane was fitted with CH4:H2O mixtures (Öberg et al., 2008).
 Currently, research on methane continues with the recent James Webb Space Telescope (JWST) data.
-The deformation mode of solid methane at 7.7 m is described by methane-bearing H2O ices: at 15 K (H2O:CH4 = 10:1111Mixtures used in fitting procedures in the cited work are shown with this notation, Rocha et al., 2024) in IRAS 2A and IRAS 23385+6053, at 13 K (H2O:CH4 = 100:6, Chen et al., 2024) in B1-c, at 12 K (H2O:CH4 = 10:1, Rocha et al., 2025) in Ced 110 IRS4, H2O and CO2 environment at 12 K222Note that the spectrum that is referred to in fitting procedure is not listed in the original paper (Rocha et al., 2017) (H2O:CO2:CH4 = 10:1:1, McClure et al., 2023) in NIR38 and J110621, using a complex mixture at 10 K (H2O:CH3OH:CO2:CH4 = 0.6:0.7:1:0.1, Yang et al., 2022) in IRAS 15398–3359 and H2O:CH4 = 10:1 at 8 K and CO2:CH4 = 5:1 at 27 K (Nakibov et al., 2025) in IRAS 23385+6053.
-According to these findings on JWST data the methane abundance in cold clouds and star-forming regions is 1.6–3.7% with respect to solid H2O.
-There are three phases of solid methane under vacuum conditions: phase I is stable above 20.4 K, phase II — below 20.4 K (e.g.
-Gerakines and Hudson, 2015b), and the third phase, stable if methane is deposited below 10 K (Colwell et al., 1962; Krupskii et al., 1973; Emtiaz et al., 2020).
+The deformation mode of solid methane at 7.7 m is described by methane-bearing H2O ices: at 15 K (H2O:CH4 = 10:1111Mixtures used in fitting procedures in the cited work are shown with this notation.; Rocha et al., 2024) in IRAS 2A and IRAS 23385+6053, at 13 K (H2O:CH4 = 100:6; Chen et al., 2024) in B1-c, at 12 K (H2O:CH4 = 10:1; Rocha et al., 2025) in Ced 110 IRS4, for an H2O and CO2 environment at 12 K222Note that the spectrum that is referred to in the fitting procedure is not listed in the original paper (Rocha et al., 2017).
+(H2O:CO2:CH4 = 10:1:1; McClure et al., 2023) in NIR38 and J110621, using a complex mixture at 10 K (H2O:CH3OH:CO2:CH4 = 0.6:0.7:1:0.1; Yang et al., 2022) in IRAS 15398–3359, and H2O:CH4 = 10:1 at 8 K and CO2:CH4 = 5:1 at 27 K (Nakibov et al., 2025) in IRAS 23385+6053.
+According to these findings, the methane abundance in cold clouds and star-forming regions is 1.6–3.7% with respect to solid H2O.
+There are three phases of solid methane under vacuum conditions: phase I is stable above 20.4 K, phase II is stable below 20.4 K (e.g., Gerakines and Hudson, 2015b), and the third phase is stable when methane is deposited below 10 K (Colwell et al., 1962; Krupskii et al., 1973; Emtiaz et al., 2020).
 Following Emtiaz et al.
-(2020) we refer to this third phase as crystalline phase II*.
-However, the literature spectra of methane typically used for interpretation of observational data are obtained at a temperature of 10 K and above, and the laboratory data on ice analogues are present as water-based mixtures for phases I and II (Hudson and Moore, 1999; Ehrenfreund et al., 1999; Rocha et al., 2017; Mifsud et al., 2023).
-Binary mixtures of methane with H2O, CO, CH3OH, NH3, CO2, N2, O2 for the first time were described in Boogert et al.
-(1997) at 10 K yet the spectral data needed for fitting is unavailable.
-There are also reflection-mode studies (e.g.
-Kaiser et al., 2014; Maity et al., 2014; Förstel et al., 2017; Bergantini et al., 2018) that characterize methane photochemistry using relevant laboratory ices, however, their spectra are not directly comparable to astronomical observations of ices in transmission.
-In this paper we present laboratory spectra of binary methane-containing ices in transmission regime, that are currently missing in the literature.
-These mixtures correspond to the environment of the polar layer of ice on the dust grains and aimed for interpretation of JWST data.
-The ices have astrochemically relevant component ratios: CH4:H2O = 1:10, CH4:CO2 = 1:5 and 1:15, CH4:CH3OH = 1:3, CH4:NH3 = 1:3.
-The set of mixtures is obtained at a substrate temperature of 10 K.
-However, in cold prestellar cores temperature may reach as low as 6 K, which is confirmed by, e.g.
-observations of gas and dust in centres of isolated cold dense cores L183 and L1544 (Pagani et al., 2004, 2007; Crapsi et al., 2007; Pagani et al., 2015).
-Also, physical modeling of L1544 resulted in the dust temperature in the core center equal to 6.5 K (Keto and Caselli, 2010).
-The presence of dust colder than 10 K is also confirmed by observations obtained with SCUBA at 850 m (Evans et al., 2001; Pattle et al., 2015).
-Therefore the set was extended with ices grown at 6.7 K with same composition.
-The extended set explores the difference between the ices grown at 6.7 K and 10 K substrate temperatures, which correspond to crystalline II* and crystalline II phases of pure methane.
-The transmission spectra of methane in binary mixtures grown at 6.7 K and CH4:CH3OH, CH4:NH3 at 10 K are presented here for the first time.
-The band strengths of methane in various environment are provided as well.
-In our previous paper (Nakibov et al., 2025) we described gaseous and solid methane in IRAS 23385+6053 using some of the spectra presented here.
-In this paper we show the utility of our new infrared spectra by exploring the spatial distribution of methane in JWST spectra of the variable protostar B335-IRS (hereafter B335).
-B335 is an isolated nebula at a distance of 90–120 pc with a young class 0 protostar IRAS19347+0727 (age yr, Yıldız et al., 2015).
-It is associated with protostellar high-velocity Herbig-Haro (HH) jet (Reipurth et al., 1992; Gâlfalk and Olofsson, 2007).
-This source has a high level of accretion and two outflows (Bjerkeli et al., 2019) and had a luminosity increase by a factor of 5–7 (Evans et al., 2023).
-This source is a natural laboratory to study ice chemistry at the early stages of star formation.
-While the sublimation of complex ices and the rich gas-phase chemistry near the heated core was observed at millimeter wavelengths (Lee et al., 2025), we focus here on the solid methane features at 7.7 m located within the outflow cavity wall.
+(2020), we refer to this third phase as the crystalline phase II*.
+However, the literature spectra of methane that were typically used to interpret observational data were obtained at temperatures of 10 K and above, and the laboratory data on ice analogs are presented as water-based mixtures for phases I and II (Hudson and Moore, 1999; Ehrenfreund et al., 1999; Rocha et al., 2017; Mifsud et al., 2023).
+Binary mixtures of methane with H2O, CO, CH3OH, NH3, CO2, and N2, O2 were described for the first time by Boogert et al.
+(1997) at 10 K, but the spectral data needed for fitting are not available.
+There are also reflection-mode studies (e.g., Kaiser et al., 2014; Maity et al., 2014; Förstel et al., 2017; Bergantini et al., 2018) that characterized methane photochemistry using relevant laboratory ices, but their spectra are not directly comparable to astronomical observations of ices in transmission.
+We present laboratory spectra of binary methane-containing ices in the transmission regime that are currently missing in the literature.
+These mixtures correspond to the environment of the polar layer of ice on the dust grains and are meant to help us interpret JWST data.
+The ices have astrochemically relevant component ratios: CH4:H2O = 1:10, CH4:CO2 = 1:5 and 1:15, CH4:CH3OH = 1:3, and CH4:NH3 = 1:3.
+The set of mixtures was obtained at a substrate temperature of 10 K.
+However, in cold pre-stellar cores, temperatures can reach as low as 6 K, which was confirmed, for instance, by observations of gas and dust in the centers of the isolated cold dense cores L183 and L1544 (Pagani et al., 2004, 2007; Crapsi et al., 2007; Pagani et al., 2015).
+Moreover, the physical modeling of L1544 yielded a dust temperature in the core center equal to 6.5 K (Keto and Caselli, 2010).
+The presence of dust colder than 10 K was also confirmed by observations obtained with Submillimetre Common-User Bolometer Array (SCUBA) at 850 m (Evans et al., 2001; Pattle et al., 2015).
+Therefore, the set was extended with ices grown at 6.7 K with the same composition.
+The extended set explored the difference between the ices grown at 6.7 K and 10 K substrate temperatures, which correspond to the crystalline II* and crystalline II phases of pure methane.
+The transmission spectra of methane in binary mixtures grown at 6.7 K and CH4:CH3OH and CH4:NH3 at 10 K are presented here for the first time.
+The band strengths of methane in various environments are provided as well.
+In our previous paper (Nakibov et al., 2025), we described gaseous and solid methane in IRAS 23385+6053 using some of the spectra presented here.
+In this paper, we show the utility of our new IR spectra by exploring the spatial distribution of methane in JWST spectra of the variable protostar B335-IRS (hereafter B335).
+B335 is an isolated nebula at a distance of 90–120 pc with a young class 0 protostar, IRAS19347+0727 (age yr; Yıldız et al., 2015).
+It is associated with a protostellar high-velocity Herbig-Haro jet (Reipurth et al., 1992; Gâlfalk and Olofsson, 2007).
+This source has a high level of accretion and two outflows (Bjerkeli et al., 2019), and its luminosity increased by a factor of 5–7 on Modified Julian Date (MJD) 56948 (Evans et al., 2023).
+This source is a natural laboratory for studying ice chemistry in the early stages of star formation.
+While the sublimation of complex ices and the rich gas-phase chemistry near the heated core were observed at millimeter wavelengths (Lee et al., 2025), we focus here on the solid methane features at 7.7 m that are located within the outflow cavity wall.
 This allows us to probe the composition of ices in the cold envelope of a young protostar.
 2 Experimental methods
-All experiments on growing analogues of interstellar ice were carried out on Ice Spectroscopy Experimental Aggregate (ISEAge) that utilizes ultra-high vacuum setup.
+All experiments on growing analogs of interstellar ice were carried out on the Ice Spectroscopy Experimental Aggregate (ISEAge), which uses an ultrahigh vacuum setup.
 A detailed description of our installation is provided in Ozhiganov et al.
 (2024).
-Briefly, the molecules are deposited on the Ge substrate from two sides via ‘background deposition’.
-The 6 mm diameter circle area (S=0.28 cm2) of the window is available for the deposition of the ice.
-The substrate is mounted on a cryogenic finger of a closed cycle He cryostat, and the whole chamber is continuously evacuated to the base pressure of × mbar.
-The operating temperatures are within the 6.7–305 K range with an error 0.1 K.
-The setup is equipped with a Stanford Research Systems RGA200 quadrupole mass spectrometer and Thermo Scientific Nicolet iS50 Fourier Transform Infrared (FTIR) spectrometer operating within a 4000–630 cm-1 (2.5–15.9 m) with 1 cm-1 resolution.
+Briefly, the molecules are deposited on the Ge substrate from two sides via background deposition.
+A circular area of the window with a diameter of 6 mm (S=0.28 cm2) is available for the deposition of the ice.
+The substrate is mounted on a cryogenic finger of a closed-cycle He cryostat, and the whole chamber is continuously evacuated to the base pressure of × mbar.
+The operating temperatures are within the 6.7–305 K range with an error of 0.1 K.
+The setup is equipped with a Stanford Research Systems RGA200 quadrupole mass spectrometer and a Thermo Scientific Nicolet iS50 Fourier transform infrared (FTIR) spectrometer operating within 4000–630 cm-1 (2.5–15.9 m) with 1 cm-1 resolution.
 The points in the collected spectra are spaced at 0.12 cm-1 intervals due to the zero filling performed on the interferometer level.
-The infrared spectra are collected in transmission regime.
+The IR spectra are collected in the transmission regime.
 The FTIR spectrometer and optical path outside the main chamber are flushed by pure N2.
 Gaseous species are introduced into the main chamber via two independent leak valves.
-We followed the same experimental procedure for growing of all ices.
+We followed the same experimental procedure to grow all ices.
 The component ratio was controlled during the deposition by means of the quadrupole mass spectrometer (QMS).
-In all experiments both components of a binary mixture were deposited simultaneously.
-The first component introduced into the main chamber was the pure methane and the second component was a matrix molecule introduced shortly after the desired ion current of methane was reached.
+In all experiments, the two components of a binary mixture were deposited simultaneously.
+The first component introduced into the main chamber was the pure methane, and the second component was a matrix molecule introduced shortly after the desired ion current of methane was reached.
 The mixture deposition time started when ion currents for both components reached the desired values.
-This approach allows to correct the influence of ion interference and overlapping ion currents, and maintain the same ion current for methane across all the experiments.
-The ion current values were taken from the calibration curves obtained for our setup via the procedure similar to Slavicinska et al.
+This approach allowed us to correct for the effect of ion interference and overlapping ion currents and to maintain the same ion current for methane in all the experiments.
+The ion current values were taken from the calibration curves obtained for our setup in a procedure similar to that described by Slavicinska et al.
 (2023).
 Briefly, the rate at which gaseous molecules are adsorbed on the substrate depends nonlinearly on the ion current and is unique to each molecule.
-By growing pure ices at different ion currents we can obtain this dependency via experiment allowing for a more precise control over the components ratio of mixed ices.
-The rates for the calibration curves are determined by measuring the absorbance on the IR spectrum output using the literature band strength values.
-Since the depositions were conducted on the Ge substrate cooled to 10 K and 6.7 K we used the band strength for methane in phase II (corrected for density) from Bouilloud et al.
-(2015) and for methane in phase II* from Gerakines and Hudson (2015b), correspondingly.
-For both sets of experiments the methane deposition rate was fixed at cm-2 s-1 and the second component deposition rate varied based on the desired components ratio.
-Deposition continued for 120 minutes.
-Thus, we get a column density of cm-2 for CH4.
-This approach enables the comparison of methane deformation (1300 cm-1) and stretching (3000 cm-1) modes in various molecular environment.
+By growing pure ices at different ion currents, we obtain this dependence via an experiment, which gives us more precise control over the component ratio of mixed ices.
+The rates for the calibration curves were determined by measuring the absorbance on the IR spectrum output using the literature band strength values.
+Since the depositions were conducted on the Ge substrate cooled to 10 K and 6.7 K, we used the band strength for methane in phase II (corrected for density) from Bouilloud et al.
+(2015) and for methane in phase II* from Gerakines and Hudson (2015b).
+For both sets of experiments, the methane deposition rate was fixed at cm-2 s-1, and the second component deposition rate varied based on the desired ratio of the components.
+The deposition continued for 120 minutes.
+We obtained a column density of cm-2 for CH4.
+This approach enabled us to compare the methane deformation (1300 cm-1) and stretching (3000 cm-1) modes in various molecular environments.
 The IR spectra were recorded every 45 seconds (an average of 32 scans).
-For the analysis in each experiment the last recorded spectrum of deposition was used to avoid the secondary adsorption effect after closing the leak valve.
-This effect is present when working with volatile species such as CH4, CO or N2 both in static conditions and during the warm up, see Appendix A for details.
-For calculations of peak positions and areas spectra are baseline-corrected for overall slope and local thin-film interference artifacts.
-Repeated measurements (50 iterations) with randomized local baseline adjustments are analyzed to minimize noise effects.
-The final values are derived from the averaged results.
-The compounds used are as follows: CH4 (99.999 %, Ugra-PGS), CO2 (99.9999 %, Ugra-PGS), NH3 (99.9 %, UralKrioGaz), deionized H2O, CH3OH (99.8 %, Vekton).
-Gaseous CO2 and CH4 are introduced directly into the dosing lines from the commercially acquired gas bottles.
-H2O, NH3 and CH3OH were purified by a freeze-pump-thaw cycle three times before each experiment.
-Before each deposition every compound is checked for purity by means of mass-spectrometry.
+For the analysis, we used the last recorded deposition spectrum in each experiment to avoid the secondary adsorption effect after closing the leak valve.
+This effect is present when working with volatile species such as CH4, CO, and N2 in both static conditions and during the warm-up (see Appendix A for details).
+To calculate the peak positions and areas, the spectra were baseline-corrected for the overall slope and local thin-film interference artifacts.
+Repeated measurements (50 iterations) with randomized local baseline adjustments were analyzed to minimize noise effects.
+The final values were derived from the averaged results.
+The compounds we used were CH4 (99.999 %, Ugra-PGS), CO2 (99.9999 %, Ugra-PGS), NH3 (99.9 %, UralKrioGaz), and deionized H2O and CH3OH (99.8 %, Vekton).
+Gaseous CO2 and CH4 were introduced directly into the dosing lines from the commercially acquired gas bottles.
+H2O, NH3, and CH3OH were purified via a freeze-pump-thaw cycle three times before each experiment.
+Before each deposition, every compound was checked for purity by means of mass spectrometry.
 3 Results
 3.1 Binary mixtures with CH4
-In this section we present a series of experiments on obtaining the binary ices of methane with molecules that constitute the polar layer of the interstellar ice mantles.
-We obtained binary mixtures at a temperature of 10 K, with methane molecules embedded in a matrix of H2O, CO2, CH3OH or NH3.
-The component ratio was chosen to qualitatively reflect current astrochemical view on species abundances in interstellar ices in cold dense molecular clouds (Boogert et al., 2015; Goto et al., 2021; McClure et al., 2023; Jiménez-Serra et al., 2025).
+In this section we present a series of experiments we conducted to obtain the binary ices of methane with molecules that constitute the polar layer of the interstellar ice mantles.
+We obtained binary mixtures at a temperature of 10 K, with methane molecules embedded in a matrix of H2O, CO2, CH3OH, or NH3.
+The component ratio was chosen to qualitatively reflect the current astrochemical view on species abundances in interstellar ices in cold dense molecular clouds (Boogert et al., 2015; Goto et al., 2021; McClure et al., 2023; Jiménez-Serra et al., 2025).
 Additionally, the comparative set was obtained at the minimum stable temperature of the ISEAge setup at 6.7 K.
-Methane absorption bands in mixtures at 6.7 K spectra differ both from pure methane and 10 K mixtures features.
-We present the spectra of mixtures for both temperatures in Fig.
+Methane absorption bands in mixtures at 6.7 K spectra differ from both pure methane and 10 K mixture features.
+We present the spectra of the mixtures for both temperatures in Fig.
 1.
 Additional experiments concerning pure methane and its transition between phases II* and II are presented in Appendix A.
-The methane absorption bands change shape and position in mixed ices, the components ratio and peak position of 3000 cm-1 stretching and 1300 cm-1 deformation mode are presented in Table 1.
+The methane absorption bands change shape and position in mixed ices.
+The component ratio and peak position of the 3000 cm-1 stretching and 1300 cm-1 deformation mode are presented in Table 1.
 The band strengths for the deformation mode of methane in various mixtures are also included.
-Methane deformation mode in CH4:CO2 = 1:5 mixture stands out because of its structure.
-It displays an unresolved double peak both in 6.7 K and 10 K experiments, which becomes resolved on heating after 30–40 K (Boogert et al., 1997; Nakibov et al., 2025).
-Moreover, this mixture displays the most distinction between the two temperatures in both vibrational modes.
-This is caused by pure CO2 having an amorphous and crystalline form as shown in Gerakines and Hudson (2015a).
-For our fixed column density of 2.121017 cm-2 pure CO2 fully crystallizes at 10 K and only partially at 6.7 K.
+The methane deformation mode in the CH4:CO2 = 1:5 mixture stands out because of its structure.
+It displays an unresolved double peak in both the 6.7 K and 10 K experiments, which becomes resolved by heating after 30–40 K (Boogert et al., 1997; Nakibov et al., 2025).
+Moreover, this mixture displays the strongest distinction between the two temperatures in both vibrational modes.
+This is caused by the amorphous and crystalline form of pure CO2, as shown by Gerakines and Hudson (2015a).
+For our fixed column density of 2.121017 cm-2, pure CO2 fully crystallizes at 10 K and crystallizes only partially at 6.7 K.
 An addition of methane, however, prevents CO2 from crystallizing at 6.7 K.
-To further investigate this we obtained more diluted mixtures: CH4:CO2 = 1:15 at 6.7 K and 10 K, the differences persisted despite the higher CO2 concentration.
-We report, therefore, the peak shape of methane differs significantly in crystalline and amorphous CO2.
-We also note that the double peak gets resolved at 10 K in CH4:CO2 = 1:15 mixture, which is in line with Hudgins et al.
-(1993), where the double structure is observed in CH4:CO2 = 1:20 mixture at 10 K.
-Methane mixtures with CH3OH and NH3 show similar features in 1300 cm-1 region and have close band strengths values.
+To further investigate this, we obtained more diluted mixtures: CH4:CO2 = 1:15 at 6.7 K and 10 K.
+The differences persisted despite the higher CO2 concentration.
+We therefore report that the peak shape of methane differs significantly in crystalline and amorphous CO2.
+We also note that the double peak becomes resolved at 10 K in the CH4:CO2 = 1:15 mixture, which is in line with the results of Hudgins et al.
+(1993), who observed the double structure in a CH4:CO2 = 1:20 mixture at 10 K.
+Methane mixtures with CH3OH and NH3 show similar features in the 1300 cm-1 region and have similar band strength values.
 The features are redshifted by 2 cm-1 from the H2O:CH4 feature.
-This opens a possibility of singling out methane not related to water ice in observational spectra, an idea first introduced by Boogert et al.
+This opens the possibility of singling out methane that is not related to water ice in observational spectra, an idea first introduced by Boogert et al.
 (1996).
-This is important, as normally methane is assumed to be formed simultaneously with water and ammonia ice via hydrogenation of carbon atoms at the onset of the formation of icy grain mantles (Qasim et al., 2020; Lamberts et al., 2022; Ferrero et al., 2023).
-However, methane can also be formed as a product of photoprocessing or ion-bombardment of methanol-rich ices (Islam et al., 2014) at later stages of ice evolution or via electron-induced radiolysis (Sullivan et al., 2016).
-The presented data allows potentially to explore the relative importance of both scenarios of methane formation.
+This is important, as normally methane is assumed to be formed simultaneously with water and ammonia ice via the hydrogenation of carbon atoms at the onset of the formation of icy grain mantles (Qasim et al., 2020; Lamberts et al., 2022; Ferrero et al., 2023).
+However, methane can also be formed as a product of the photo-processing or ion-bombardment of methanol-rich ices (Islam et al., 2014) in later stages of the ice evolution or via electron-induced radiolysis (Sullivan et al., 2016).
+The presented data could potentially be used to explore the relative importance of the two scenarios of methane formation.
 | Mixture | Component ratio | Temperature (K) | Peak position (cm-1) | Peak position (cm-1) | Band Strength (cm) |
-| CH4 | - | 8 | 3010.0 | 1296.9 | 1.04×10-17aaaaThe band strength is taken from Gerakines and Hudson (2015b); |
+| CH4 | - | 8 | 3010.0 | 1296.9 | 1.04×10-17aaaaThe band strength is taken from Gerakines and Hudson (2015b). |
 | CH4 | - | 10 | 3008.8 | 1301.3 | 8.4×10-18bbbbThe band strength is taken from Bouilloud et al. (2015). |
@@
 3.2 B335 methane survey
-To test the usability of the obtained spectra we fitted the 7.7 m methane absorption band of the Class 0 protostar B335.
+To test the usability of the obtained spectra, we fitted the 7.7 m methane absorption band of the Class 0 protostar B335.
 JWST Mid-IR Instrument (MIRI) Medium Resolution Spectrograph (MRS) observations toward B335 are provided within the Investigating Protostellar Accretion proposal (PID 1802, P.I.
@@
 Megeath; Megeath et al., 2021).
-Level 3 observation data were taken from the MAST database (MAST DOI: 10.17909/yr06-tf19), we further scaled the CH2-SHORT channel data to match the CH1-LONG signal level.
-MIRI MRS observations included three levels of JWST calibration pipeline with version 1.19.1 and Calibration References Data System (CRDS) context jwst_1413.pmap (CRDS_VER = ’12.1.11’) describing in Greenfield and Miller (2016); Bushouse et al.
-(2024); van Gelder et al.
+Level 3 observation data were taken from the Mikulski Archive for Space Telescopes (MAST) database (DOI: 10.17909/yr06-tf19), and we further scaled the CH2-SHORT channel data to match the CH1-LONG signal level.
+MIRI MRS observations included three levels of the JWST calibration pipeline with version 1.19.1 and Calibration References Data System (CRDS) context jwst_1413.pmap (CRDS_VER = ’12.1.11’), which is described in Greenfield and Miller (2016), Bushouse et al.
+(2024), and van Gelder et al.
 (2024).
-Unfortunately, the overlap region coincides with the OCN- absorption band and we excluded this specie from the further analysis.
-B335 is viewed edge-on and has two prominent outflows (e.g.
-Bjerkeli et al., 2019), the west of which displaying strong continuum emission around 7.7 m.
+Unfortunately, the overlap region coincides with the OCN- absorption band, and we excluded this species from further analysis.
+The object B335 is viewed edge-on and has two prominent outflows (e.g., Bjerkeli et al., 2019), the western of which displays strong continuum emission around 7.7 m.
 Based on the intensity map at 8 m (Fig.
-2) we picked multiple apertures that cover bright pixels.
-This approach maximizes the signal-to-noise ratio, enhancing the analysis of the finer spectral details.
-To date, this is the first study that explores the spatial distribution on MIRI-MRS JWST data for a molecule with relatively low abundance relative to H2O.
-We note that a spatial distribution study was carried out recently for solid H2O, CO and CO2 on the NIRCam data (Smith et al., 2025).
-The spectra were extracted from the cylindrical apertures: seven 1.53′′ apertures were used to assess the spatial distribution of methane and a 4.59′′ aperture was used to ensure consistency between fits in smaller apertures.
+2), we selected multiple apertures that covered bright pixels.
+This approach maximizes the signal-to-noise ratio, which facilitates our analysis of the finer spectral details.
+This is the first study that explores the spatial distribution on MIRI-MRS JWST data for a molecule with a low abundance relative to H2O.
+We note that a spatial distribution study was carried out recently for solid H2O, CO, and CO2 on NIRCam data (Smith et al., 2025).
+The spectra were extracted from the cylindrical apertures: seven 1.53′′ apertures were used to assess the spatial distribution of methane, and a 4.59′′ aperture was used to ensure consistency between fits in smaller apertures.
 The center positions for the chosen apertures are listed in Table 2.
 All spectra were processed in the same way.
-After baseline subtraction a linear combination of methane laboratory mixtures was fitted to the observational spectrum.
-The linear scale coefficients were obtained by minimizing , given as:
+After baseline subtraction, a linear combination of methane laboratory mixtures was fitted to the observational spectrum.
+The linear scale coefficients were obtained by minimizing , given as
 where and are observational and modeled optical depths, respectively, and is the point-wise standard deviation of the observational spectrum.
-The standard deviation was obtained from the residuals between the original optical depth spectrum and a smoothed curve obtained with locally weighted scatterplot smoothing (LOWESS, Cleveland, 1979) with a 24-point window.
-Column densities of H2O and CO2 were estimated based on 6 m and 15.2 m bands, respectively.
-The fitted spectra are shown in the Fig.
+The standard deviation was obtained from the residuals between the original optical depth spectrum and a smoothed curve obtained with locally weighted scatterplot smoothing (Cleveland, 1979) with a 24-point window.
+The column densities of H2O and CO2 were estimated based on the 6 m and 15.2 m bands, respectively.
+The fitted spectra are shown in Fig.
 3.
@@
 2.
-Fitting results show that about 70% and 30% of methane budget in B335 are best described by methane in H2O and methane in CO2 environments, respectively (see Table 2).
-A minor part of methane is found in CH3OH environment, however, the detection at this level may be sensitive to e.g.
-baseline selection.
-Spectrum is excluded from the analysis due to low signal-to-noise ratio.
-The 1.53′′ apertures display variations in ice composition with CH4 in CO2 to CH4 in H2O relative abundances decreasing towards the protostar.
-There is also an overall increase in column density towards the core, likely explained by an increase of density towards the center.
-The presence of CH4 in CO2 matrix provides some insights on ice formation at early stages of the protostar.
+The fitting results show that about 70% and 30% of the methane budget in B335 are best described by methane in H2O and methane in CO2 environments, respectively (see Table 2).
+A minor fraction of the methane is found in the CH3OH environment, but the detection at this level may be sensitive to baseline selection, for example.
+Spectrum was excluded from the analysis because the signal-to-noise ratio was low.
+The 1.53′′ apertures display variations in ice composition, with CH4 in CO2 to CH4 in H2O relative abundances decreasing toward the protostar.
+The column density toward the core also increases overall, which is likely explained by an increase in density toward the center.
+The presence of CH4 in the CO2 matrix provides some insights into ice formation in early stages of the protostar.
 An interesting correlation is observed between CH4 and CO2.
-The Spearman’s and p-values (Spearman, 1904) for CO2 and CH4 column densities (, =0.007) and abundances (, =0.003) indicate strong, statistically significant, positive correlation between these species.
-The correlations between CH4 and CO2 are known from as early as Spitzer methane survey (Öberg et al., 2008).
-Modeling presented in Garrod and Pauly (2011) shows that under certain conditions methane and CO2 can form concurrently, occupying the innermost layers of ice mantles.
+The Spearman and p values (Spearman, 1904) for the CO2 and CH4 column densities (, =0.007) and abundances (, =0.003) indicate a strong statistically significant positive correlation between these species.
+The correlations between CH4 and CO2 are known from the Spitzer methane survey (Öberg et al., 2008).
+The modeling presented in Garrod and Pauly (2011) shows that under certain conditions, methane and CO2 can form concurrently, occupying the innermost layers of ice mantles.
 Ferrero et al.
-(2024) explored the reactivity of 3C-OH2 complex with various species, revealing barrierless reactions with atomic O and H.
-These reactions lead to production of CH2 and CO, which are the direct precursors to CH4 and CO2, respectively.
-While parallel formation of methane and CO2 is not largely explored, these results can be interpreted as evidence towards the concurrence for these processes.
+(2024) explored the reactivity of the 3C-OH2 complex with various species, revealing barrierless reactions with atomic O and H.
+These reactions lead to the production of CH2 and CO, which are the direct precursors of CH4 and CO2, respectively.
+While the parallel formation of methane and CO2 has not been largely explored, these results can be interpreted as evidence that these processes occur.
 | Aperture | Coordinates | (CH4) | CH4:H2O | CH4:CO2 | CH4:CH3OH | (CO2) | (H2O) | (CH4)/(H2O) |
@@
 4 Astrochemical implications and conclusions
-In this paper we present spectra of pure methane and binary mixtures with methane for 6.7 K and 10 K deposition temperatures.
-The following mixtures are presented in this paper: CH4:H2O = 1:10, CH4:CO2 = 1:5, CH4:CO2 = 1:15, CH4:CH3OH = 1:3 and CH4:NH3 = 1:3.
+We have presented the spectra of pure methane and binary mixtures with methane for 6.7 K and 10 K deposition temperatures: CH4:H2O = 1:10, CH4:CO2 = 1:5, CH4:CO2 = 1:15, CH4:CH3OH = 1:3, and CH4:NH3 = 1:3.
 All spectra of methane in astrochemically relevant mixture ratios are presented for the first time at 6.7 K.
-The new band strength values and spectra obtained at temperature below 9 K reveal the following astrochemical implications:
+The new band strength values and spectra obtained at a temperature below 9 K revealed the astrochemical implications listed below.
 1.
-Currently, the band strength of pure methane at 10 K is used majorly for column density estimations in observational data (e.g.
-McClure et al., 2023; Rocha et al., 2024, 2025).
+The band strength of pure methane at 10 K is currently mainly used for column density estimations in observational data (e.g., McClure et al., 2023; Rocha et al., 2024, 2025).
 However, solid interstellar methane is found in mixtures with other icy species rather than in pure form.
-We also note that interstellar ices differ from laboratory analogues both in how they are produced and how they are processed.
+We also note that interstellar ices differ from laboratory analogs both in how they are produced and in how they are processed.
 Thus, while the band strengths in Table 1 provide a better basis for interpreting observations, they are still approximate.
-Nevertheless, our new data invites a discussion on recent literature estimates.
-Since the methane band strengths in mixtures are 20 % higher on average than the band strength of pure methane at 10 K, the column densities of interstellar methane may be overestimated by this value in e.g.
-McClure et al.
-(2023).
-Interestingly, this partially explains the underestimation of solid methane abundance by astrochemical models toward background stars NIR38 and J110621 reported in Jiménez-Serra et al.
+Nevertheless, our new data invite a discussion on recent literature estimates.
+Since the methane band strengths in mixtures are 20 % higher on average than the band strength of pure methane at 10 K, the column densities of interstellar methane may be overestimated by this value by McClure et al.
+(2023), for instance.
+Interestingly, this partially explains the underestimation of the solid methane abundance by astrochemical models toward the background stars NIR38 and J110621 reported in Jiménez-Serra et al.
 (2025).
 2.
-We present the the first publicly available digital spectral library of methane absorbance spectra in non-H2O environment, providing the possibility to identify non-H2O methane environments on newly obtained JWST spectra.
-Solid methane is typically considered to reside in polar water-rich environment due its early formation along with water ice (see e.g.
-Hasegawa et al., 1992; Clément et al., 2023; Borshcheva et al., 2025).
-This is further supported by successful fits of methane features in JWST spectra of interstellar ices with CH4:H2O mixtures (e.g.
-Chen et al., 2024; Rocha et al., 2024, 2025).
-However, there are possible evidences for its presence in mixtures with other molecules.
-For example, works by Boogert et al.
+We have presented the the first publicly available digital spectral library of methane absorbance spectra in a non-H2O environment, providing the possibility to identify non-H2O methane environments in newly obtained JWST spectra.
+Solid methane is typically considered to reside in a polar water-rich environment because it forms early along with water ice (see, e.g., Hasegawa et al., 1992; Clément et al., 2023; Borshcheva et al., 2025).
+This is further supported by successful fits of methane features in JWST spectra of interstellar ices with CH4:H2O mixtures (e.g., Chen et al., 2024; Rocha et al., 2024, 2025).
+However, there is possible evidence that it is present in mixtures with other molecules.
+For example, Boogert et al.
 (1996) and Nakibov et al.
-(2025) show that observed shapes of 7.7 m methane feature can be partially interpreted with non-H2O methane-containing mixtures.
+(2025) show that the observed shapes of the 7.7 m methane feature can be partially interpreted with non-H2O methane-containing mixtures.
 Additionally, icy methane mixed with apolar species such as N2 and CO is found on icy bodies in the outer Solar System (e.g., Emtiaz et al., 2022).
 The obtained spectra of ices of astrochemical ratios are published on Zenodo (DOI: 10.5281/zenodo.15350566).
-Comparison between spectra obtained at 6.7 K and 10 K reveal temperature-induced changes in shape and peak position, although we find it unlikely that these differences are prominent enough to be discerned within the observational data.
-However, it is known that central regions of prestellar cores are well-shielded from external heating radiation.
-Consequently, as shown both observationally (Pagani et al., 2004, 2007; Crapsi et al., 2007; Pagani et al., 2015) and theoretically (Keto and Field, 2005; Keto and Caselli, 2010, and references therein), dust temperature there may reach 6 K.
-According to a recent modeling by Borshcheva et al.
-(2025), approximately 80% of solid methane in the prestellar core L1544 resides on grains with temperature below 9 K.
-Thus, to accurately interpret future observations of ices on a line of sight toward L1544, laboratory spectra and band strengths of mixtures obtained both at 6.7 K and 10 K might be needed.
-The presented spectra can be directly compared to observational data, because the deformation mode of methane is unlikely to be affected by grain size, depending mostly on temperature and molecular environment (Boogert et al., 1997).
+Our comparison of spectra obtained at 6.7 K and 10 K revealed temperature-induced changes in shape and peak position, although we find it unlikely that these differences are prominent enough to be discerned within the observational data.
+However, it is known that the central regions of pre-stellar cores are well shielded from external heating radiation.
+Consequently, as shown both observationally (Pagani et al., 2004, 2007; Crapsi et al., 2007; Pagani et al., 2015) and theoretically (Keto and Field, 2005; Keto and Caselli, 2010, and references therein), the dust temperature there can reach 6 K.
+According to recent modeling by Borshcheva et al.
+(2025), approximately 80% of solid methane in the pre-stellar core L1544 resides on grains with temperatures below 9 K.
+Thus, to accurately interpret future observations of ices on a line of sight toward L1544, laboratory spectra and band strengths of mixtures obtained at both 6.7 K and 10 K might be needed.
+The presented spectra can be directly compared to observational data because the deformation mode of methane is unlikely to be affected by grain size; it mostly depends on temperature and the molecular environment (Boogert et al., 1997).
 3.
-We present the first spatial survey performed on JWST MIRI-MRS data, exploring methane in B335.
-The results show that about 30% of solid methane is found in CO2 environment, rather than in H2O matrix.
-The consistency between findings in full and smaller apertures proves the usability of presented spectra for JWST data analysis.
-The observed ice composition gradient further highlights that JWST provides enough sensitivity for such studies, prompting further research.
-5 Data availability
-The laboratory infrared absorbance spectra in transmission regime in the 2.5–15.9 m (4000–630 cm-1) region for CH4 in different environments at 6.7 K and 10 K can be found on Zenodo at the following link: https://zenodo.org/records/15350566.
+We have presented the first spatial survey performed on JWST MIRI-MRS data and explored methane in B335.
+The results show that about 30% of solid methane is found in a CO2 environment and not in the H2O matrix.
+The consistency between the findings from full and smaller apertures proves that the spectra we have presented can be used to analyze JWST data.
+The observed ice composition gradient further demonstrates that JWST provides enough sensitivity for such studies, which should motivate further research.
+Data availability
+The laboratory IR absorbance spectra in the transmission regime in the 2.5–15.9 m (4000–630 cm-1) region for CH4 in different environments at 6.7 K and 10 K can be found on Zenodo via the following link: https://zenodo.org/records/15350566.
 Acknowledgements.
@@
 Katerina Borshcheva for providing the information on the distribution of methane ice in the model of L1544.
-We thank the anonymous reviewer for their comments that helped us to improve the manuscript.
+We thank the anonymous reviewer for their comments that helped us to improve the manuscript and A&A Language Editor Astrid Peter for corrections in the English text.
 This research work is funded by the Russian Science Foundation via 23-12-00315 agreement.References
@@
 There are three phases of solid methane under vacuum conditions.
-The crystalline phase II is stable below 20.4 K and has the fcc lattice with eight sublattices: six of which have site symmetry D2d and two — site symmetry Oh (Kobashi et al.
+The crystalline phase II is stable below 20.4 K and has the fcc lattice with eight sub-lattices: six of which have site symmetry D2d and two of which have site symmetry Oh (Kobashi et al.
 1977).
@@
 Phase II* cannot be obtained by cooling methane ice from the triple point or from phase II ice.
-On warm-up phase II* ice undergoes irreversible transition to phase II.
+On warming phase II* ice undergoes irreversible transition to phase II.
 The second transition temperature depends on experimental conditions due to metastability and is observed below 10 K (Colwell et al.
@@
 2020).
-Despite the extensive research carried out on methane solid phases in infrared at low temperatures in recent decades (Kobashi et al.
+Despite the extensive research carried out on methane solid phases in the IR at low temperatures in recent decades (Kobashi et al.
 1977; Hudgins et al.
@@
 Gerakines and Hudson (2015b) report that the majority of their spectra were obtained in the 8–12 K range, but the spectra are referred to as obtained ‘at 10 K’ for simplicity throughout the text.
-Thus, it is not clear at what temperature exactly were published experiments carried out, a point that also applies to Hudson et al.
+Thus, it is not clear at what temperature exactly the published experiments were carried out, a point that also applies to Hudson et al.
 (2015).
@@
 1973).
-We attempted to obtain the spectrum at their Fig.
-1(f) at precisely 10 K, but failed even using a lower deposition rate than stated in their paper.
+We attempted to obtain the spectrum in their Fig.
+1f at precisely 10 K, but failed even using a lower deposition rate than stated in their paper.
 Therefore, prior to obtaining the set of methane-bearing mixtures at 10 K we decided to investigate the transition between phase II* and phase II.
 Under the deposition rate and duration of experiments described in the main text methane retained phase II* structure when deposited at 6.7 K throughout the whole deposition time.
-In contrast, for deposition at 10 K, the double peak structure characteristic of phase II started to form once ice reached certain column density, see Figure 4.
+In contrast, for deposition at 10 K, the double peak structure characteristic of phase II started to form once the ice reached certain column density; see Fig.
+4.
 At a deposition rate of 5.9 cm-2 s-1 the transition started after 35 minutes, corresponding to a column density of 1.21016 cm-2.
@@
 The change in peak structure occurs at a critical value of 1.21016 cm-2.
-This equals approximately to 6 monolayers of pure methane ice on each side of the Ge substrate, given that single monolayer column density is 11015 cm-2.
+This equals approximately to 6 monolayers of pure methane ice on each side of the Ge substrate, given that a single monolayer column density is 11015 cm-2.
 From our experiments we conclude that phase II* is stable at 6.7 K.
@@
 A definitive answer would come from X-ray diffractometry study as performed in Mizuno et al.
-(2016, 2022), which showed that other molecular ices that are ‘conventionally amorphous’ for astrochemistry can exhibit positionally ordered but orientationally disordered structure (CO or N2), while other (CO2, CS2) are indeed amorphous.
+(2016, 2022), which showed that other molecular ices that are ‘conventionally amorphous’ for astrochemistry can exhibit positionally ordered but orientationally disordered structure (CO or N2), while others (CO2, CS2) are indeed amorphous.
 We also note an instrumental effect that concerns both temperature programmed desorption (TPD) and static temperature experiments with volatile species.
@@
 (2025).
-The evacuation therefore is governed by the amount of temporarily stored methane rather than by turbomolecular pump’s efficiency, resulting in a ‘tail’ in QMS signal after deposition.
+The evacuation therefore is governed by the amount of temporarily stored methane rather than by the turbomolecular pump’s efficiency, resulting in a ‘tail’ in the QMS signal after deposition.
 During TPD warm surfaces release stored material early, polluting the QMS signal, as in Collings et al.
 (2004).
-Released gas can get readsorbed by the cold substrate.
-An example of that can be found in the cyclic experiment in Gerakines and Hudson (2015b) (see their Fig.
+Released gas can get re-adsorbed by the cold substrate.
+An example of that can be found in the cyclic experiment in Gerakines and Hudson (2015b, see their Fig.
 1, right column).
 We note that the integrated intensity increased after the first heating-holding-cooling cycle, indicating an increase in column density.
-Second cycle did not cause an increase as stored methane depleted on first heating.
+The second cycle did not cause an increase because the stored methane was depleted on the first heating.
 This increase did not affect the results of their paper.